Calculate the area bounded by the 
y = sinx and x-axis from 0 to 2pi.

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

f(x) = sinx
Area ‘A’
A = \int_{0}^{\pi }sinxdx + |\int_{\pi }^{2\pi }sinxdx|
A = (-cosx)_{0}^{\pi }+ |\int_{\pi }^{2\pi }sinxdx|
A = 2+ |\int_{\pi }^{2\pi }sinxdx|
A = 2+ |(-cosx)_{\pi }^{2\pi }|
A = 2 + |-2|
A = 4